OpenStudy (anonymous):

Proof d/dx e^x=e^x using limit def (1+ 1/h)^h h->infinity somehow e limit definition implies d/dx e^x=e^x

OpenStudy (anonymous):

\[\lim_{h \rightarrow 0} (e ^{x+h} - e ^{x}) / h\]

OpenStudy (anonymous):

That is the definition of the derivative of e^x, and you have to proof that that equals e^x First of all, we have to on it a bit: \[(e ^{x+h} - e ^{x})/h = e ^{x}(e ^{h}-1)/h\] ` But, then, the first limit is equal to \[[e ^{x}\lim_{h \rightarrow 0} (e ^{h}-1)/h\] But, that limit, \[\lim_{h \rightarrow 0} (e ^{h}-1)/h = 1\] So \[e ^{x}\lim_{h \rightarrow 0} (e ^{h}-1) = e ^{x}.1 = e^x\]

OpenStudy (anonymous):

I was able to figure it out ^^ thanks

OpenStudy (anonymous):

But we haven´t proof yet that \[\lim_{h \rightarrow 0}(e^h-1) = 1\], so, we have to proof it. First of all you have to keep in mind the following substitution: Let t, \[t = 1/(e^h-1) \] , then, \[1/t = e^h-1 \] and, adding 1 at both sides \[(1/t)+1 = e^h \] and then,taking ln at both sides you get....

OpenStudy (anonymous):

you get \[\ln (1/t +1) = h\] Then, think for a moment about the limit of that if h approaches 0...then ln(1/t +1) has to approach 0....And ln becomes 0 when it argument becomes 1. So, if h approaches to 0, ln has to approach 0 too, and 1/t + 1 has to approach 1, but the only way is if t approaches infinity. So, if h approaches 0, then t approaches infinity. Then we are able to proof the limit...

OpenStudy (anonymous):

\[\lim_{h \rightarrow o} (e^h - 1)/h =\lim_{t \rightarrow \infty}(1/t)/\ln(1+1/t) \] So, at this step i have only applied the substitution t that we have written before, and note that at right side we have the limit of t approaching infinity...because the variable change and the consideration about the limits that I did at the last post. After this point, you have only to work on the expression, and you´ll arrive to the definition of the e number that the exercise suggested \[\lim_{t \rightarrow \infty}(1/t)/\ln(1+1/t) = \lim_{t \rightarrow \infty} 1/\ln(1+1/t)^{1/t} = \lim_{t \rightarrow \infty} \ln (e) = 1\] At the final step I just used the definition of e number \[\lim_{t \rightarrow \infty} (1+1/t)^t = e\] Well, thats all...=)

OpenStudy (anonymous):

Oh sorry, I haven´t read you!! Haha, no problem man! Congratulations then

OpenStudy (anonymous):

thanks though haha